# RingMoE: Mixture-of-Modality-Experts Multi-Modal Foundation Models for Universal Remote Sensing Image Interpretation

### v1

@@ 4.3.1 Generalization Analysis of Remote Sensing Foundation Models
To evaluate the generalization capability of RingMoE’s learned representations, we freeze its backbone parameters and fine-tune only the task-specific heads across three downstream tasks: scene classification, object detection, and semantic segmentation. As shown in Fig.13.a, recent foundation models [66, 68, 69, 17, 27] suffer significant performance degradation under this setting. For instance, SatLas [68] exhibits a 28.98% accuracy drop on the AID scene classification dataset, indicating limited generalization of its pre-trained features across diverse RS tasks. In contrast, RingMoE maintains high performance across all three tasks, achieving 95.82% on AID classification, 74.90% on DIOR detection, and 67.39% on iSAID segmentation. These results demonstrate that RingMoE effectively captures robust and transferable feature representations during pretraining, enabling strong performance even with a frozen backbone.
@@ 4.3.2 Discussion of RingMoE’s Usage for Downstream
To enhance RingMoE’s resource efficiency while maintaining accuracy, we evaluate its performance using parameter-efficient fine-tuning (PEFT). Specifically, the Convpass strategy [85] introduces lightweight adaptation modules into the pre-trained RingMoE, freezing the core model weights while fine-tuning only these modules. Fig.13.b compares three fine-tuning strategies, i.e., freezing, PEFT, and full fine-tuning. Remarkably, RingMoE with PEFT achieves near-equivalent accuracy to full fine-tuning for tasks such as semantic segmentation, while using only 7.5% of the parameters (i.e., 80M) required for full fine-tuning. This balance between precision and reduced resource demands underscores RingMoE’s suitability for real-world remote sensing applications, enabling efficient and rapid interpretation on resource-constrained systems.
@@ 4.3.3 Effectiveness of Multi-modal Pre-training
4.3.4 Necessity and Impact of Extending Foundation model to 10 Billion-scale Parameters
The practical importance of developing a 10-billion-scale parameter foundation model merits close investigation. To explore this, we pre-train a 1-billion-parameter (1B) Swin Transformer v2 using SimMIM self-supervised method on the same optical datasets as RingMoE. As illustrated in Fig.13.d, the 14.7-billion-parameter (14.7B) RingMoE model (allocated as 6.5B parameters per modality) demonstrates substantially higher performance than the 1B Swin Transformer v2. Remarkably, even after parameter compression to 1 billion (RingMoE-KC), RingMoE maintains its advantage over the Swin Transformer v2 baseline. These results highlight the significant capabilities of the 14.7B RingMoE model in advancing remote sensing tasks. Additionally, by offering a 1-billion parameter variant, RingMoE extends its accessibility to resource-limited applications, thereby broadening its impact and usability in practical settings.
@@ 5 Conclusion
This paper presents RingMoE, a 14.7-billion-parameter multi-modal foundation model that achieves SOTA performance on 23 out of 25 public benchmarks across six key RS tasks. Pre-trained on 400 million multi-modal images, including Opt, MS, and SAR data, RingMoE demonstrates strong generalization and robust few-shot performance, making it particularly effective in data-scarce scenarios. In addition to accuracy, RingMoE’s modular design allows for flexible adaptation to diverse downstream tasks through expert pruning, maintaining competitive performance even in compressed versions. These results underscore the potential of large-scale multi-modal foundation models in RS, with promising applications in disaster response, environmental monitoring, and urban planning. Future work could further enhance its capabilities by incorporating LiDAR and hyper-spectral data, leveraging sensor-specific properties for more targeted self-supervised learning.
@@ Acknowledgments
  • RingMoE implementation (Appendix B): This section provides a detailed discussion of the design and core components of the RingMoE model. It covers the collection and pre-processing of the pre-training dataset, the architectural details of the model, and the specifics of the pre-training process. Furthermore, it details the implementation of RingMoE for downstream tasks, including dataset descriptions, experimental setups, and evaluation metrics.
@@ B.1 Collection and Pre-processing of RingMoE Pre-training Dataset
The RingMOSS pre-training dataset is designed to provide a comprehensive and multi-dimensional analysis of RS data, incorporating diverse data sources, including optical (Opt), multi-spectral (MS), and Synthetic Aperture Radar (SAR) imagery (see Fig.A1-A2). Opt data is sourced from high-resolution aerial and satellite platforms, such as JL-1, GF-2, GeoEye, WorldView, QuickBird, IKONOS, and SPOT, with spatial resolutions ranging from 0.3 meters to 30 meters. MS images, capturing four bands—blue, green, red, and near-infrared—are obtained from the GF-1 and GF-2 satellites. The GF-1 PMS camera provides images at a resolution of 8 meters, while the GF-2 PMS camera delivers higher-resolution images at 3.2 meters. SAR data is sourced from the GF-3 satellite, encompassing both SAR-L1 (complex-valued form) and SAR-L2 (amplitude form) images. The SAR-L1 images, with a resolution of 1 meter, include four polarization modes—HH, HV, VH, and VV—combined into eight-channel images for pre-training. These images capture both amplitude and phase information, providing high-resolution insights into Earth’s surface. SAR-L2 images, derived from SAR-L1 data, undergo additional processing, including geo-coding and radiometric calibration.
The RingMOSS dataset is constructed by aggregating RS images from both public and proprietary sources, with a focus on ensuring cross-regional coverage, temporal variability, and diversity in scenes and objects. It spans urban, agricultural, forest, water, and mountainous regions, thereby providing broad representational coverage. During the pre-processing stage, all original RS images are partitioned into 192$\times$192 pixel tiles using a sliding window technique with a step size of 64 pixels. Due to edge effects, sensor artifacts, and data stitching issues, black borders often appear along the image edges. To prevent these artifacts from affecting model training, any tile containing more than 30% black border is discarded, leaving only valid regions for further processing and training. Additionally, regions affected by cloud cover are automatically detected and removed during this stage. These steps ensure the high quality and integrity of the dataset, providing reliable input for subsequent model training and analysis.
@@ B.3 Implementation for Pre-training
For pre-training, all multi-modal images adopt the following pre-processing settings, including randomly cropping the image in the range of 0.2$\times$-1.0$\times$ of the original size, resizing to 192$\times$192, applying horizontal flipping and normalization operations. For Opt, MS, and SAR-L2 images, they are input as 8-bit data formats with 3 channels, 4 channels, and 1 channel, respectively. For SAR-L1 images, which consist of four polarization modes and eight channels, we maintain the 16-bit format to maximize the preservation of complex signal information.
@@
B.4 Implementation for Downstream Tasks
@@ B.4.1 Scene Classification
  • NWPU-RESISC45 (Opt) [60], is a publicly available remote sensing image scene classification (RESISC) dataset containing 31,500 images of size 256$\times$256, with ground sample distances (GSD) ranging from 0.2 meters to 30 meters. It covers 45 diverse scene classes including “Aircraft”, “Airports”, “Baseball diamond”, “Basketball court”, “Beach”, “Bridge”, “Jungle”, “Church”, “Circular farmland”, “Cloud”, “Commercial area”, “Dense housing”, “Desert”, etc., with 700 images in each category. Following previous work [16, 27], we allocate 10% of the dataset for training and reserve the remaining 90% for testing.
  • AID (Opt) [65], is a large-scale dataset of aerial imagery sourced from Google Earth, containing 10,000 images of size 600$\times$600, with GSD ranging from 0.5 meters to 8 meters. This dataset spans 30 distinct aerial scene classes, including “Airport”, “Bare ground”, “Baseball stadium”, “Beach”, “Bridge”, “Center”, “Church”, “Business”, “Dense housing”, etc., with 220 to 400 images in each category. Following previous work [16, 27], we allocate 20% of the dataset for training and reserve the remaining 80% for testing.
(2) Implementation Detail. Our experiments are performed in the mmpretrain framework [^1] https://github.com/open-mmlab/mmpretrain. RingMoE is employed to extract features while a linear mapping layer is employed for classification. The optimizer utilizes AdamW with an initial learning rate of 5e-5 and a weight decay of 0.05. The learning rate scheduler consists of two phases: a linear scheduler for the first 10 epochs with an initial factor of 0.01, followed by a cosine scheduler starting from the 10th epoch with a minimum learning rate of 5e-6. The training process continues for a total of 200 epochs. For NWPU-RESISC45, the images are randomly cropped to 224$\times$224 and for AID, the images are randomly cropped to 512$\times$512. To enhance model robustness, we implement data augmentation techniques including RandomCrop, RandomFlip, and RandomErasing for both datasets.
@@ B.4.2 Semantic Segmentation
  • iSAID (Opt) [63], is an instance-level semantic segmentation dataset consisting of 2806 aerial images captured by various satellite sensors. The Image resolutions range from 800$\times$800 to 4000$\times$13000. It includes 655,451 instances across 15 object categories, with non-object pixels labeled as background, totaling 16 categories. The dataset is split into training, validation, and testing sets, comprising 1411, 458, and 937 samples. Following [16, 27], we evaluate the model performance on the validation set.
  • ISPRS Potsdam (Opt), is a scene-level semantic segmentation dataset that contains 38 high-resolution aerial images with a ground sampling distance (GSD) of 0.05 meters, each with a fixed size of 6000$\times$6000 pixels. Following the setup in RingMo [16], we conduct experiments using images that include near-infrared, red, and green spectral bands, focusing on five categories: “Impervious surface”, “Building”, “Low vegetation”, “Tree”, and “Car”. The dataset is split into 24 images for training and 14 for testing.
  • Dyna.-pla (MS) [71], is a multi-spectral scene-level semantic segmentation dataset captured from the PlanetFusion satellite. It includes images from 75 different locations, with each featuring 24 images taken at different times, accompanied by annotations for 7 land use and land cover semantic classes. Each image is captured in four bands (Red, Green, Blue, and Near-Infrared), with a ground sampling distance (GSD) of 3 meters and an image size of 1024$\times$1024. Following the official leaderboard configuration, the dataset is divided into 55 locations for training, 10 for validation, and 10 for testing. In our experiments, we utilize the official validation and test sets for performance evaluation.
  • SARSegL1 (SAR-L1) [72], is a fully polarized semantic segmentation dataset, captured by the GF3 satellite over Guangzhou and Hangzhou, China, at a spatial resolution of 5 meters. It includes two images with dimensions of 5456$\times$4708 and 6192$\times$4888 pixels, respectively, with each image comprising four polarization modes: HH, HV, VH, and VV. Each mode has two channels that capture amplitude and phase information. The dataset is annotated with five categories: “Water”, “Vegetation”, “Bare ground”, “Building”, and “Road” annotated by professionals. To facilitate model training and evaluation, the images are cropped into 500$\times$500 pixels with an overlap of 200, resulting in a total of 540 samples, with an 8:2 split between training and testing sets.
  • Air-POLSAR-SEG (SAR-L2) [73], is a scene-level semantic segmentation dataset, contains 500 single-channel SAR images (in amplitude form) with a spatial resolution of 8 meters. Each 512$\times$512 pixel image provides pixel-level annotations across six categories: “Housing area”, “Industrial area”, “Natural area”, “Land-use area”, “Water”, and “Other area”. The dataset is divided into training and test sets in a 7:3 ratio, enabling precise land-use classification tasks in remote sensing applications. Due to the small proportion of Land-use and Other areas, these two categories are excluded from the experimental results.
(2) Implementation Detail. All semantic segmentation downstream tasks are implemented using the mmsegmentation framework [^2] https://github.com/open-mmlab/mmsegmentation. We adopt Upernet [86] as the segmentation framework and utilize compressed RingMoE-KC as the backbone for feature extraction. The optimizer is AdamW with an initial learning rate of 1e-4 and a weight decay of 0.05. The learning rate schedule consists of two phases: a linear warm-up for the first 1,500 iterations with a factor of 0.1, followed by a polynomial scheduler with a minimum learning rate of 1e-6. For the iSAID dataset, images are cropped to 896$\times$896, and the model is trained for 80k iterations with a batch size of 8. For the Potsdam and Air-POLSAR-SEG datasets, images are cropped to 512$\times$512, with training also set at 80k iterations and a batch size of 8. For the Dyna.-pla dataset, images are resized to 1024$\times$1024, with training set at 16k iterations and a batch size of 8. For the SARSegL1 dataset, the two-channel data from four polarization modes are concatenated to create an 8-channel sample, with images cropped to 500$\times$500 and trained for 20k iterations with a batch size of 4. In our experiments, we employ regular data augmentation strategies, including RandomFlip, RandomScaling (from 0.5 to 2.0), and RandomCrop.
@@
B.4.3 Multi-modal Semantic Segmentation

### v2

@@ 4.3.1 Generalization Analysis of Remote Sensing Foundation Models
To evaluate the generalization capability of RingMoE’s learned representations, we freeze its backbone parameters and fine-tune only the task-specific heads across three downstream tasks: scene classification, object detection, and semantic segmentation. As shown in Fig.12.a, recent RSFMs [66, 68, 69, 16, 26] suffer significant performance degradation under this setting. For instance, SatLas [68] exhibits a 28.98% accuracy drop on the AID classification dataset, indicating limited generalization of its pre-trained features across diverse RS tasks. In contrast, RingMoE maintains high performance across all three tasks, achieving 95.82% on AID classification, 74.90% on DIOR detection, and 67.39% on iSAID segmentation. These results demonstrate that RingMoE effectively captures robust and transferable feature representations during pretraining, enabling strong performance even with a frozen backbone.
@@ 4.3.2 Discussion of RingMoE’s Usage for Downstream
To enhance RingMoE’s resource efficiency while maintaining accuracy, we evaluate its performance using parameter-efficient fine-tuning (PEFT). Specifically, the Convpass strategy [90] introduces lightweight adaptation modules into the pre-trained RingMoE, freezing the core model weights while fine-tuning only these modules. Fig.12.b compares three fine-tuning strategies, i.e., freezing, PEFT, and full fine-tuning. Remarkably, RingMoE with PEFT achieves near-equivalent accuracy to full fine-tuning for tasks such as semantic segmentation, while using only 7.5% of the parameters (i.e., 80M) required for full fine-tuning. This balance between precision and reduced resource demands underscores RingMoE’s suitability for real-world RS applications, enabling efficient and rapid interpretation on resource-constrained systems.
TABLE V: Performance comparison between original RingMoE@Swin-1B and a lightweight variant (RingMoE@Swin-88M), evaluated on two RS classification datasets. Both models use identical settings; the only difference lies in the backbone size (88M vs. 1B).
Method Backbone
NWPU-RESISC45
(TR=10%)
AID
(TR=20%)
OA(%) OA(%)
RingMoE RMoE@Swin-88M 93.64 96.57
RMoE@Swin-1B 95.90 98.19
TABLE VI: Ablation study of the Expert Design of RMoE on four key RS tasks, including optical classification, SAR detection, multi-spectral segmentation, and multi-modal segmentation. We evaluate the impact of each expert type by selectively removing them.
Expert Design in RMoE Optical Cls. SAR-L2 Det. Multi-spectral Seg. Multi-modal Seg.
NWPU-RESISC45 [58] HRSID [80] Dyna.-pla [72] WHU-OPT-SAR [75]
Modal-specialized Collaborative Shared Acc(%) mAP50(%) mIoU(%) mIoU(%)
94.15 93.66 47.09 52.63
93.78 93.24 46.28 53.39
94.59 93.92 47.37 54.33
95.05 94.20 47.60 54.70
@@ 4.3.3 Effectiveness of Multi-modal Pre-training
4.3.4 Effectiveness of Scaling Foundation Models to 10 Billion-scale Parameters
To investigate the effectiveness of scaling RSFMs to the 10B-parameter level, we conduct a series of ablation studies. Specifically, we pre-train a 1B-parameter Swin Transformer v2 baseline (without the RMoE layer) using the SimMIM [7] self-supervised framework on the same optical datasets as RingMoE. As illustrated in Fig.12.d, the full 14.7B RingMoE model (with 6.5B parameters allocated per modality) consistently outperforms the 1B baseline across multiple downstream tasks, demonstrating the advantages of scaling through sparse expert modeling. Furthermore, even the 1B dense variant, RingMoE-KC, surpasses the original Swin baseline, highlighting the effectiveness of expert pruning strategy in retaining high-capacity knowledge.
To further validate the impact of parameter scaling, we replace the FFN layers in the Swin-B (88M) backbone with RMoE layers, forming a lightweight variant (RingMoE@Swin-88M), and compare it against the original RingMoE@Swin-1B. Both models share identical settings, differing only in backbone size (88M vs. 1B). As reported in Tab.V, RingMoE@Swin-1B consistently outperforms RingMoE@Swin-88M on the NWPU-RESISC45 [58] and AID [65] datasets, providing rigorous evidence that increasing parameters alone yields significant performance gains.In summary, these findings confirm the necessity and effectiveness of parameter scaling for enhancing the representational capacity of RSFMs.
4.3.5 Ablation Studies on Model Components
Expert Design. We conduct ablation experiments by selectively removing each expert type from RingMoE. As reported in Tab.VI, removing modal-specialized experts leads to notable performance drops in single-modal tasks (-1.27% accuracy on optical classification, -1.32% mIoU on multi-spectral segmentation), highlighting their effectiveness in capturing modal-specific knowledge. Excluding collaborative experts leads to a 2.07% mIoU decline on multi-modal segmentation, underscoring their role in facilitating cross-modal representation learning. Removing the shared expert consistently causes minor performance declines across all tasks, indicating its stabilizing role in maintaining globally transferable features. These results collectively demonstrate the complementary benefits of the three expert types and underscore the necessity of our hybrid expert design for robust multimodal learning.
Physics-informed Loss Function. To evaluate the impact of incorporating SAR physical priors, we replace the proposed physics-informed reconstruction loss with a pixel-wise reconstruction loss during SAR-L1 pre-training. As reported in Appendix Tab.18, this substitution leads to consistent performance degradation on both SAR-L1 and SAR-L2 downstream tasks (–5.67% mIoU on SARSegL1, –3.14% mAP on SAR-AIRCraft). Notably, although the physics-informed loss is applied only to SAR-L1 data during pretraining, performance gains are also observed on SAR-L2 tasks. This improvement stems from collaborative experts in RMoE enabling implicit cross-modal knowledge transfer. These findings indicate that embedding physics-based constraints effectively enhances both modal-specific representation learning and generalization across SAR branches.
5 Limitations and Future Work
Despite the strong performance of RingMoE, several limitations remain. First, the pre-training corpus consists of unpaired multi-modal RS imagery, limiting the use of cross-modal self-supervised paradigms such as contrastive learning. Second, the RingMOSS dataset suffers from regional and modality imbalance, with data predominantly sourced from a few countries. Third, while expert pruning enables efficient deployment, pre-training the full-scale RingMoE remains computationally intensive. Fourth, the benefits of multi-modal fusion are only verified on two benchmarks due to the scarcity of geo-aligned multi-modal RS datasets. To address these limitations, we plan to curate large-scale, geographically diverse, and geo-registered multi-modal RS datasets to support fine-grained cross-modal learning and enhance generalization. We will also explore further model compression via knowledge distillation to enable deployment on resource-limited platforms (e.g., UAVs). Moreover, RingMoE will be extended as a universal vision encoder for RS multi-modal large language models and integrated with prompt-based frameworks such as SAM to enhance interactivity and generality in referential perception tasks.
@@ 6 Conclusion
This paper presents RingMoE, a 14.7-billion-parameter multi-modal foundation model that achieves SOTA performance on 23 out of 25 public benchmarks across six key RS tasks. Pre-trained on 400 million multi-modal images, including Opt, MS, and SAR data, RingMoE demonstrates strong generalization and robust few-shot performance, making it particularly effective in data-scarce scenarios. In addition to accuracy, RingMoE’s modular design allows for flexible adaptation to diverse downstream tasks through expert pruning, maintaining competitive performance even in compressed versions. These results underscore the potential of large-scale multi-modal foundation models in RS, with promising applications in disaster response, environmental monitoring, and urban planning.
@@ Acknowledgments
  • RingMoE implementation (Appendix B): This section provides a detailed discussion of the design and core components of the RingMoE model. It covers the collection and pre-processing of the pre-training dataset, the architectural details of the model, the specifics of the pre-training process, and the implementation of expert pruning. Furthermore, it details the implementation of RingMoE for downstream tasks, including dataset descriptions, experimental setups, and evaluation metrics.
@@
  • Comparison with popular MLLMs (Appendix D): This section compares the proposed RingMoE model with recent Multimodal Large Language Models (MLLMs) in RS scenarios. The evaluation focuses on task specialization, modality generalization, and performance.
  • More comprehensive ablation studies and analysis (Appendix E): This section conducts additional ablation analyses, covering the comparison of expert pruning strategies, the effect of SAR-L1 pretraining on improving SAR understanding, and the impact of the physics-informed loss applied during SAR-L1 pretraining.
@@ B.1 Collection and Pre-processing of RingMoE Pre-training Dataset
The RingMOSS pre-training dataset is designed to provide a comprehensive and multi-dimensional analysis of RS data, incorporating diverse data sources, including optical (Opt), multi-spectral (MS), and Synthetic Aperture Radar (SAR) imagery (see Fig.A1-A2). Opt data is sourced from high-resolution aerial and satellite platforms, such as JL-1, GF-2, GeoEye, WorldView, QuickBird, IKONOS, and SPOT, with spatial resolutions ranging from 0.3 meters to 30 meters. MS images, capturing four bands—blue, green, red, and near-infrared—are obtained from the GF-1 and GF-2 satellites. The GF-1 PMS camera provides images at a resolution of 8 meters, while the GF-2 PMS camera delivers higher-resolution images at 3.2 meters. SAR data is sourced from the GF-3 satellite, encompassing both SAR-L1 (complex-valued form) and SAR-L2 (amplitude form) images. The SAR-L1 images, with a resolution of 1 meter, include four polarization modes, HH, HV, VH, and VV, combined into eight-channel images for pre-training. These images capture both amplitude and phase information, providing high-resolution insights into Earth’s surface. SAR-L2 images, derived from SAR-L1 data, undergo additional processing, including geo-coding and radiometric calibration.
The RingMOSS dataset is constructed by aggregating RS images from both public and proprietary sources, with a focus on ensuring cross-regional coverage, temporal variability, and diversity in scenes and objects. It spans urban, agricultural, forest, water, and mountainous regions, thereby providing broad representational coverage. During the pre-processing stage, all original RS images are partitioned into $\times192$ pixel tiles using a sliding window technique with a step size of 64 pixels. Due to edge effects, sensor artifacts, and data stitching issues, black borders often appear along the image edges. To prevent these artifacts from affecting model training, any tile containing more than 30% black border is discarded, leaving only valid regions for further processing and training. Additionally, regions affected by cloud cover are automatically detected and removed during this stage. These steps ensure the high quality and integrity of the dataset, providing reliable input for subsequent model training and analysis.
Due to licensing restrictions and confidentiality agreements, we are unable to release data from private sources. However, to support reproducibility and community research, we have open-sourced a curated subset of RingMOSS [^2] All public data, along with detailed documentation are available at: https://github.com/HanboBizl/RingMoEDatasets. that is fully derived from public sources. This released subset retains all four modalities and follows the exact same preprocessing and formatting protocols as the original dataset used for pretraining.
@@ B.3 Implementation for Pre-training
[…]
@@
B.4 Implementation for Expert Pruning
B.4.1 Sparse Expert Pruning Schemes
Although we introduce a load-balancing loss (Eq.(4) in the main text) to encourage uniform expert activation, we consistently observe an uneven distribution of routing decisions across cooperative experts, often influenced by the input modality. Specifically, different modalities tend to activate distinct subsets of cooperative experts in each RMoE layer, leading to modality-dependent expert preferences despite the balancing regularization. To quantify this phenomenon, we compute the modal-specific activation frequency of each cooperative expert $E_{k}^{C}$ for modality $m$, formally defined as:
$f(E_{k}^{C}|x_{m})=\frac{1}{N_{m}}\sum_{i=1}^{N_{m}}\mathbb{I}\left[G^{C}(x_{m;i})=E_{k}^{C}\right]$ (15)
Where $N_{m}$ denotes the total number of tokens from modality $m$, and $x_{m;i}$ is the $i^{th}$ token. The indicator function $\mathbb{I}(\cdot)$ equals 1 if expert $E_{k}^{C}$ is selected for the token $x_{m;i}$, and 0 otherwise. Here, $k\in{1,2,…,N_{C}}$ indexes the cooperative experts, with $N_{C}$ representing the total number of cooperative experts in each RMoE layer. This frequency metric captures the degree to which each expert is utilized when processing data from different modalities.
We consistently observe a long-tailed distribution: a small number of experts dominate the routing decisions for each modality, while many others remain underutilized. Moreover, these preferences show consistent modality-specific patterns:
  • Optical and multi-spectral modalities tend to activate similar sets of experts, reflecting their shared spectral characteristics.
  • SAR-L1 and SAR-L2 also show close alignment in expert selection, though SAR-L1 activates a broader range of experts, potentially due to its richer spatial structure.
  • In contrast, optical and SAR modalities activate almost disjoint expert subsets, which is consistent with their fundamentally different imaging mechanisms (e.g., passive reflection vs. active microwave emission).
These observations motivate a modality-aware, layer-wise sparse expert pruning strategy, designed to eliminate those cooperative experts that contribute little to the representation learning of any modality. For each RMoE layer and modality m, we rank all cooperative experts by their activation frequency $f(E_{k}^{C}|x_{m})$, then retain only the top 25% most frequently activated ones. The pruning threshold $\varphi$ is determined as the 75th percentile of the activation frequency distribution:
$\varphi=\text{Percentile}_{75}\left(\left\{f(E_{k}^{C}|x_{m})\right\}_{k=1}^{N_{C}}\right)$ (16)
Experts with $f(E_{k}^{C}|x_{m})<\varphi$ are deemed underutilized and are pruned for modality m. After computing $\varphi$ for all modalities, we take the union of the retained experts across modalities at each layer. This ensures that experts important to any single modality are preserved, while those consistently underused are pruned.
We validate the pruning design on the NPWU classification dataset using our 6.5B-parameter RingMoE model. As shown in Tab.A1, we compare both fixed-threshold and percentile-based strategies under various retention ratios. The Top-25% scheme, which retains the most frequently activated 25% of cooperative experts for each modality and layer, achieves the best trade-off between compression and performance, reducing model size from 6.5B to 4.3B parameters with only a 0.47% drop in classification accuracy.
TABLE A1: Ablation study on the pruning threshold $\varphi$ selection. We evaluate both fixed and percentile-based pruning strategies using the 6.5B RingMoE-Opt (optical modality) model on the NPWU classification dataset.
Selection Strategy #Params Acc(%)
Fixed $\varphi=0.1$ 4.7B 95.06
Top 25% 4.3B 95.43
Top 50% 5.0B 95.57
Top 75% 5.7B 95.71
No Pruning 6.5B 95.90
B.4.2 Dense Expert Integration Schemes
Knowledge Compressing. Beyond the two aforementioned schemes of direct expert knowledge aggregation, we aim to preserve key knowledge from each expert while ensuring effective compression. Low-rank decomposition is a powerful technique for model compression, enabling the transformation of a high-dimensional matrix into a low-rank representation while maximally retaining its information. To achieve this, we compress each expert’s weight matrix into a low-rank form [47] and subsequently merge them into a high-rank matrix with the same dimensionality as the original, as illustrated in Fig.7 of the main text.
$W_{1}=M_{1}^{S}+M_{1}^{C}+W_{1}^{Shared}$ (20)
B.5 Implementation for Downstream Tasks
@@ B.5.1 Scene Classification
  • NWPU-RESISC45 (Opt) [58], is a publicly available remote sensing image scene classification (RESISC) dataset containing 31,500 images of size $\times256$, with ground sample distances (GSD) ranging from 0.2 meters to 30 meters. It covers 45 diverse scene classes including “Aircraft”, “Airports”, “Baseball diamond”, “Basketball court”, “Beach”, “Bridge”, “Jungle”, “Church”, “Circular farmland”, “Cloud”, “Commercial area”, “Dense housing”, “Desert”, etc., with 700 images in each category. Following previous work [15, 92], we allocate 10% of the dataset for training and reserve the remaining 90% for testing.
  • AID (Opt) [65], is a large-scale dataset of aerial imagery sourced from Google Earth, containing 10,000 images of size $\times600$, with GSD ranging from 0.5 meters to 8 meters. This dataset spans 30 distinct aerial scene classes, including “Airport”, “Bare ground”, “Baseball stadium”, “Beach”, “Bridge”, “Center”, “Church”, “Business”, “Dense housing”, etc., with 220 to 400 images in each category. Following previous work [15, 92], we allocate 20% of the dataset for training and reserve the remaining 80% for testing.
(2) Implementation Detail. Our experiments are performed in the mmpretrain framework [^1] https://github.com/open-mmlab/mmpretrain. RingMoE is employed to extract features while a linear mapping layer is employed for classification. The optimizer utilizes AdamW with an initial learning rate of 5e-5 and a weight decay of 0.05. The learning rate scheduler consists of two phases: a linear scheduler for the first 10 epochs with an initial factor of 0.01, followed by a cosine scheduler starting from the 10th epoch with a minimum learning rate of 5e-6. The training process continues for a total of 200 epochs. For NWPU-RESISC45, the images are randomly cropped to $\times224$ and for AID, the images are randomly cropped to $\times512$. To enhance model robustness, we implement data augmentation techniques including RandomCrop, RandomFlip, and RandomErasing for both datasets.
@@ B.5.2 Semantic Segmentation
  • ISPRS Potsdam (Opt) [71], is a scene-level semantic segmentation dataset that contains 38 high-resolution aerial images with a ground sampling distance (GSD) of 0.05 meters, each with a fixed size of $\times6000$ pixels. Following the setup in RingMo [15], we conduct experiments using images that include near-infrared, red, and green spectral bands, focusing on five categories: “Impervious surface”, “Building”, “Low vegetation”, “Tree”, and “Car”. The dataset is split into 24 images for training and 14 for testing.
  • Dyna.-pla (MS) [72], is a multi-spectral scene-level semantic segmentation dataset captured from the PlanetFusion satellite. It includes images from 75 different locations, with each featuring 24 images taken at different times, accompanied by annotations for 7 land use and land cover semantic classes. Each image is captured in four bands (Red, Green, Blue, and Near-Infrared), with a ground sampling distance (GSD) of 3 meters and an image size of $\times1024$. Following the official leaderboard configuration, the dataset is divided into 55 locations for training, 10 for validation, and 10 for testing. In our experiments, we utilize the official validation and test sets for performance evaluation.
[…]
  • Air-POLSAR-SEG (SAR-L2) [74], is a scene-level semantic segmentation dataset, contains 500 single-channel SAR images (in amplitude form) with a spatial resolution of 8 meters. Each $\times512$ pixel image provides pixel-level annotations across six categories: “Housing area”, “Industrial area”, “Natural area”, “Land-use area”, “Water”, and “Other area”. The dataset is divided into training and test sets in a 7:3 ratio, enabling precise land-use classification tasks in remote sensing applications. Due to the small proportion of Land-use and Other areas, these two categories are excluded from the experimental results.
(2) Implementation Detail. All semantic segmentation downstream tasks are implemented using the mmsegmentation framework [^2] https://github.com/open-mmlab/mmsegmentation. We adopt Upernet [93] as the segmentation framework and utilize compressed RingMoE-KC as the backbone for feature extraction. The optimizer is AdamW with an initial learning rate of 1e-4 and a weight decay of 0.05. The learning rate schedule consists of two phases: a linear warm-up for the first 1,500 iterations with a factor of 0.1, followed by a polynomial scheduler with a minimum learning rate of 1e-6. For the iSAID dataset, images are cropped to $\times896$, and the model is trained for 80k iterations with a batch size of 8. For the Potsdam and Air-POLSAR-SEG datasets, images are cropped to $\times512$, with training also set at 80k iterations and a batch size of 8. For the Dyna.-pla dataset, images are resized to $\times1024$, with training set at 16k iterations and a batch size of 8. For the SARSegL1 dataset, the two-channel data from four polarization modes are concatenated to create an 8-channel sample, with images cropped to $\times500$ and trained for 20k iterations with a batch size of 4. In our experiments, we employ regular data augmentation strategies, including RandomFlip, RandomScaling (from 0.5 to 2.0), and RandomCrop.
(3) Evaluation Metric. We employ mean intersection over union (mIoU), overall accuracy (OA), mean accuracy (mAcc), and mean F1 score (mF1) to evaluate segmentation performance. mIoU calculates the segmentation accuracy per category by taking the ratio of the intersection-over-union (IoU) between predicted and true categories and then averaging across all categories to reflect the model’s capability in segmenting individual classes. For each category, the IoU can be calculated by $\text{IoU}=\frac{\text{TP}}{{\text{TP}+\text{FP}+\text{FN}}}$. Meanwhile, OA evaluates the percentage of correctly classified pixels across the dataset, providing a straightforward view of overall model accuracy. mAcc measures the average pixel accuracy for each category, defined as the ratio of correctly predicted pixels in a category to the total ground truth pixels of that category. For each category, accuracy is calculated as: $\text{Acc}=\frac{\text{TP}}{\text{TP}+\text{FN}}$. mF1 captures the harmonic mean of precision and recall for each category and averages them across all categories. For each category, the F1 score is defined as: $\text{F1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, where $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ and $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$. Together, these metrics offer a well-rounded understanding of a semantic segmentation model’s performance across different scenarios and categories.
B.5.3 Multi-modal Semantic Segmentation